# LLM as GNN: Graph Vocabulary Learning for Text-Attributed Graph Foundation Models

### v2

@@ 4.1 Graph Understanding Module
$\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\subset\mathcal{N}_{i}\right\}$ $\leftarrow\text{Prompt}_{\text{sample}}\left(X^{(l-1)},\mathcal{N}_{i}\right),$ (3)
@@
$X_{i}^{(l)}$ $\leftarrow\text{Prompt}_{\text{agg-update}}\left(\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\in\mathcal{N}_{i}^{(l)}\right\},X_{i}^{(l-1)}\right),$ (4)
@@
where we seek flexible message passing mechanism without relying on predefined operators such as mean and weighted aggregators. A discussion on handling prompt length constraints is provided in Appendix F. In terms of optimization, we refer to unsupervised graph learning with contrastive loss to learn task-agnostic representations, which encourages similarity among neighboring nodes while pushing apart distant ones. Instead of traditional negative sampling, which becomes redundant in our setting, we leverage LLM prompts to intuitively steer the contrastive objective. Details of prompt construction are provided in Appendix D.
Figure 3: The pipeline of PromptGFM. (a) Graph Understanding Module: For arbitrary TAGs from different domains, prompt-based GNN replicates traditional embedding-based GNN workflow in the text space, generating compact node representations. (b) Graph Inference Module: We establish a unified graph vocabulary and extract language-based IDs to generate massive pure-language prompts, enabling LLM fine-tuning across graphs and tasks.
@@
After repeating $L$ rounds, we obtain the final textual representation as $T_{i}=X_{i}^{(L)}$. These representations encapsulate both semantic and structural information, effectively solving the outlined issues: (a) Graph Description. One-hop neighbor descriptions are equivalent to an adjacency matrix, representing the entire graph structure. (b) Graph Structure. To capture high-order relationships, we iteratively invoke LLMs with the same prompts, feeding each round’s output into the next. (c) Graph Semantics. Simultaneously, we explicitly instruct LLMs to produce concise yet expressive textual representations for each node, gradually refining them for denser and richer semantics.
We systematically examine how prompt-based GNNs faithfully mirror embedding-based GNNs in Appendix C. In summary, our fine-grained replication exemplifies the potential of using LLMs as GNNs, fostering seamless GNN-LLM integration and elegant graph-text alignment.
4.2 Graph Inference Module
@@
The graph inference module seeks to unify diverse graphs and tasks and acquire transferable knowledge through multi-instruction fine-tuning with LLMs. Existing methods, which represent nodes as OOV tokens, suffer from semantic misalignment between ID-based and language-based embeddings in task-oriented prompts due to distinct vocabularies, constraining their transferability. To tackle this limitation, we introduce a novel language-based graph vocabulary that bridges this incompatibility, enabling the generation of large-scale, coherent, and readable instructions to support LLM inference.
@@
Graph Vocabulary. Learning a transferable graph vocabulary, whose fundamental units can represent each node, is critical for building effective GFMs. Its utility hinges on three key criteria: expressiveness, transferability, and scalability. Since each node has been equipped with a textual representation in the graph understanding module that captures its core semantics and local structure, we intuitively construct a universal language-based graph vocabulary based on these concise yet informative representations. In this vocabulary, each node is represented by a finite sequence of language tokens, known as a language-based ID, defined as:
$\mathcal{F}:V\rightarrow T,$ (5)
@@
where each graph node $v_{i}$ is assigned a sequence as ${T}_{i}=\left(t_{i}^{1},t_{i}^{2},\ldots,t_{i}^{m_{i}}\right)$, where $t_{i}^{k}\in\mathcal{X}$ and $\mathcal{X}$ is a dictionary of general natural language tokens. Collectively, the language-based IDs for all nodes in $V$ are denoted as $T=\left\{T_{0},T_{1},\dots,T_{|V|-1}\right\}$. Along this line, natural language tokens serve as the fundamental building blocks of our graph vocabulary, with structured token sequences representing each graph node, analogous to words in standard lexicons.
@@
Our vocabulary satisfies all expected criteria: (a) Expressiveness. Language-based IDs preserve rich semantic and structural information from open-world settings. (b) Transferability. Like human lexicons, the vocabulary is grounded in natural language, providing a shared foundation that ensures cross-graph interpretability and transferability. (c) Scalability. Any node, whether previously seen or not, can dynamically generate its language-based ID, maintaining compatibility with existing nodes while preventing token explosion. We analyze vocabulary coverage and uniqueness in Section 5.3.
@@
Multi-Instruction Fine-Tuning. As shown in Figure 4, we employ a multi-instruction fine-tuning framework to incorporate diverse graphs and tasks [9, 51]. We index nodes from the vocabulary and embed their language-based IDs into task-oriented templates to construct completed instructions $\mathcal{T}$:
$\mathcal{T}\leftarrow\text{Prompt}_{\text{template}}\left(T,G\mid\mathcal{F}\right),$ (6)
@@
which include language-based IDs of the central nodes, along with its corresponding local structure from $G$ depending on specific graph-centric tasks (e.g., node classification or link prediction). As these instructions are composed entirely of natural language tokens, they are fully readable and coherent. All graph-centric tasks are then reformulated in a text-to-text format, allowing the LLM to be fine-tuned within a unified instruction learning paradigm [34]. Given a target sequence $Y$, the loss function can be computed as:
$\mathcal{L}$ $=-\sum_{j=1}^{|Y|}\log\text{Pr}\left(Y_{j}\mid\mathcal{T},Y_{<j}\right),$ (7)
@@
where $\text{Pr}\left(Y_{j}\mid\mathcal{T},Y_{<j}\right)$ denotes the probability of the $j$-th token $Y_{j}$ in the output sequence $Y$, conditioned on the instruction $\mathcal{T}$ and all previous tokens $Y_{<j}=(Y_{1},Y_{2},\dots,Y_{j-1})$. To support broad generalization, we fine-tune LLM backbones such as T5 [39], FLAN [51], and Llama [45] to co-train a unified model across diverse graphs and tasks. As a result, the model acquires open-world knowledge and generalizes effectively to previously unseen graphs and tasks.
Figure 4: An instance of graph inference module in link prediction, where language-based IDs are indexed from the graph vocabulary to generate readable instructions with task-oriented templates. A multi-instruction fine-tuning framework is employed to unify graphs and tasks and learn transferable knowledge for GFMs.
@@
Constrained Decoding via Prefix Tree Search. To mitigate LLM hallucination, we introduce a constrained decoding strategy based on a prefix tree search to regulate LLM outputs  [4, 43]. Taking generative link prediction as an example, we craft a prefix tree from language-based IDs of all candidate nodes, where each tree node corresponds to a natural language token and each root-to-leaf path uniquely identifies a valid node ID. During autoregressive generation, token selection is constrained to a valid path within the prefix tree, ensuring that predictions align with actual graph nodes and eliminating hallucinations. This effectiveness can be attributed to the discrete nature of language-based IDs, further highlighting the flexibility of our proposed graph vocabulary.
5 Experiments
In this section, we conduct extensive experiments to address the following research questions (RQs):
$\bullet$ RQ1: How does PromptGFM perform on node classification and link prediction tasks?
$\bullet$ RQ2: How transferable is PromptGFM across diverse graphs and tasks as a general-purpose GFM?
$\bullet$ RQ3: How effective is the proposed graph vocabulary in terms of quantitative evaluation?
$\bullet$ RQ4: How does each module contribute to the overall performance of PromptGFM?
$\bullet$ RQ5: What is the computational efficiency in graph understanding and inference modules?
$\bullet$ RQ6: What factors influence the effectiveness of GNN replication within PromptGFM?
Datasets. We present the following datasets covering three domains: (a) computer science: Cora [32], Citeseer [13], Ogbn-arxiv [18], and WikiCS  [33]; (b) e-commerce: Photo and History [36, 57]; (c) biomedical: PubMed [42]. Details can be found in Appendix A.
Baselines. We make comprehensive comparisons with existing methods in four categories: (1) Graph-agnostic methods. We consider the MLP model without graph structure. (2) GNN-based methods. We employ four fundamental GNN models: GCN  [23], GAT [46], GraphSAGE [16], and ReVGNN [24]. We also explore SGFormer [54] and NodeFormer [53], which leverage Transformer architectures to model graph data. (3) LM-only methods. We utilize BERT [10], Sentence-BERT  [40], and Llama3-8B [45] to address downstream tasks. (4) GNN-LLM Integration methods. Following the proposed taxonomy, GraphPrompter  [29] is an instance of using GNNs to enhance LLMs. OFA  [27] and ENGINE  [62] are examples of leveraging LLMs for GNNs. Besides, we incorporate LLaGA  [6] as an attempt of implementing LLM as GNN. Details can be found in Appendix B.
[…]
@@ 5 Experiments
$p$-value 4e-3 2.73e-8 - 4.48e-10 4.85e-7 3.62e-2 9.93e-4
[…]
Method Discriminative Setting
@@
[…]
Method Generative Setting
@@
Reproduction Settings. We conduct all experiments in PyTorch on four NVIDIA RTX A6000 GPUs. The graph understanding module uses OpenAI’s GPT-4o mini while we fine-tune a Flan-T5 or Llama3-8B in the graph inference module. We employ 10-fold cross-validation and compute average results with standard deviations. For evaluation, we use accuracy and Macro-F1 (M-F1) for node classification, and accuracy and HR@1 for link prediction. To align with the TAG setting, BERT is used for node initialization across all applicable baselines. More details are provided in Appendix E.
5.1 Performance Comparison (RQ1)
Main Results. We train each model independently from scratch on a single graph and compare their performance. Tables 1 and 3 summarize the results for node classification and link prediction accuracy, respectively. PromptGFM achieves substantial improvements over state-of-the-art models, leading to the following observations. (a) Graph-based models consistently outperform graph-agnostic ones, highlighting the importance of structural information. (b) Our method surpasses OFA, ENGINE, and GraphPrompter on node classification, suggesting the limitations of their decoupled architectures in both GNN-for-LLM and LLM-for-GNN paradigms. (c) PromptGFM also surpasses LLaGA, which relies on templates to implicitly understand graph semantics and structures. These results indicate that heuristic prompting alone is insufficient to capture high-order dependencies without a true GNN mechanism. In contrast, PromptGFM effectively replicates message passing through a prompt-based GNN, demonstrating the feasibility of using LLMs as GNNs and establishing a new paradigm for graph-text alignment.
Generative Link Prediction. Following a transductive setting, we partition the graph by links, construct an input graph from the training edges, and predict missing connections for test nodes, using existing nodes as candidate targets. From Table 3, PromptGFM consistently outperforms traditional GNN models in the generative link prediction. Unfortunately, existing GNN-LLM methods fail to address this task, as their reliance on OOV token embeddings prevents LLM outputs from being directly mapped to specific nodes, often resulting in hallucinated predictions. In contrast, our work represents nodes as finite token sequences, allowing constrained decoding via prefix tree search to regulate LLM outputs. This further highlights the essence of our graph vocabulary and demonstrates the flexibility of PromptGFM in generative scenarios.
Table 4: Intra-domain cross-graph transferability
to Citeseer within the computer science domain.
@@ 5.1 Performance Comparison (RQ1)
Table 5: Inter-domain cross-graph transferability from e-commerce to biomedical domain (i.e. PubMed).
@@
5.2 Cross-graph and Cross-task Transferability (RQ2)
We investigate cross-graph transferability in both intra-domain and inter-domain scenarios. Following  [7], we explore two settings: (1) Pre-training, where models are evaluated on entirely unseen datasets without prior exposure during training, and (2) Co-training, where a portion of the target dataset is available and jointly used with auxiliary datasets for training. These two dimensions yield four distinct transfer paradigms, enabling a comprehensive assessment of cross-graph transferability.
Intra-domain Cross-graph Transferability. As illustrated in Table 5, all settings surpass direct inference using an off-the-shelf LLM. Notably, incorporating Arxiv yields greater gains than Cora, suggesting that larger and richer datasets enhance transferability more effectively. In the pre-training setting, substantial improvements over direct inference demonstrate strong zero-shot transferability within the domain. A similar, albeit smaller, trend is observed in the co-training setting, further validating the effectiveness of intra-domain knowledge transfer. These findings underscore the potential to collect more graph data and build more knowledgeable and transferable GFMs.
Table 6: Inter-domain cross-graph transferability from computer science to biomedical domain (i.e. PubMed).
@@ 5.2 Cross-graph and Cross-task Transferability (RQ2)
Inter-domain Cross-graph Transferability. We report cross-domain transfer results from e-commerce and computer science to the biomedical domain in Tables 5 and 6, respectively. In the pre-training setting, while incorporating a single external graph enhances performance, adding multiple source graphs may lead to negative transfer. More critically, in the co-training setting under supervised learning, performance consistently deteriorates as more cross-domain data is introduced. This decline stems from domain-specific knowledge conflicts, potentially caused by catastrophic forgetting or incompatible hyperparameters that fail to adapt to varying domain distributions [7, 59]. While PromptGFM offers a unified interface for diverse graphs and tasks, achieving robust and reliable cross-domain transfer remains a challenge, which we leave for future exploration.

### v3

@@ 4.1. Graph Understanding Module
(3) $\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\subset\mathcal{N}_{i}\right\}$ $\leftarrow\text{Prompt}_{\text{sample}}\left(X^{(l-1)},\mathcal{N}_{i}\right),$
@@
(4) $X_{i}^{(l)}$ $\leftarrow\text{Prompt}_{\text{agg-update}}\left(\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\in\mathcal{N}_{i}^{(l)}\right\},X_{i}^{(l-1)}\right),$
@@
where we seek flexible message passing mechanism without relying on predefined operators such as mean and weighted aggregators. A discussion on handling prompt length constraints is provided in Appendix C.1. In terms of optimization, we refer to unsupervised graph learning with a prompt-driven contrastive loss, which encourages similarity among neighboring nodes while pushing apart distant ones. As a task-agnostic approach to graph representation learning, contrastive optimization naturally aligns with the generality requirements of GFMs. Instead of traditional negative sampling, which becomes redundant in our setting, we prompt LLMs to intuitively and implicitly steer the contrastive objective. Details of prompt construction are provided in Appendix B.
@@
After repeating $L$ rounds, we obtain the final textual representation as $T_{i}=X_{i}^{(L)}$. These representations encapsulate both semantic and structural information, effectively solving the aforementioned issues: (a) Graph Description. One-hop neighbor descriptions are equivalent to an adjacency matrix, representing the entire graph structure. (b) Graph Structure. To capture high-order relationships, we iteratively invoke LLMs with the same prompts, feeding each round’s output into the next. (c) Graph Semantics. Simultaneously, we explicitly instruct LLMs to produce concise yet expressive textual representations for each node, gradually refining them for denser and richer semantics. In summary, our fine-grained replication exemplifies the potential of using LLMs as GNNs, fostering seamless GNN-LLM integration and elegant graph-text alignment.
4.2. Graph Inference Module
@@
The graph inference module seeks to unify diverse graphs and tasks and acquire transferable knowledge through multi-instruction fine-tuning with LLMs. Existing methods, which represent nodes as OOV tokens, creates semantic misalignment between ID-based and language-based embeddings in task-oriented prompts due to distinct vocabularies, and substantially limits transferability across graphs and tasks. To overcome this challenge, we introduce a novel language-based graph vocabulary that bridges this incompatibility, enabling the generation of large-scale, coherent, and readable instructions to support effective LLM inference.
4.2.1. Graph Vocabulary.
@@
Learning a transferable graph vocabulary, whose fundamental units can represent each node, is critical for building effective GFMs. Its utility hinges on three key criteria: expressiveness, transferability, and scalability. Since each node has been equipped with a textual representation in the graph understanding module that captures its core semantics and local structure, we intuitively construct a universal language-based graph vocabulary that matches LLM’s native vocabulary. In our graph vocabulary, each node is represented by a finite sequence of language tokens, known as a language-based ID, defined as:
(5) $\mathcal{F}:V\rightarrow T,$
@@
where each graph node $v_{i}$ maps a sequence ${T}_{i}=\left(t_{i}^{1},t_{i}^{2},\ldots,t_{i}^{m_{i}}\right)$, and $t_{i}^{k}\in\mathcal{X}$ and $\mathcal{X}$ is a dictionary of general natural language tokens. Collectively, the language-based IDs for all nodes in $V$ are denoted as $T=\left\{T_{0},T_{1},\dots,T_{|V|-1}\right\}$. Along this line, the structured token sequence of a graph node is analogous to a word in standard lexicons, ensuring semantic fidelity and eliminating OOV issue.
@@
Our vocabulary satisfies all expected criteria: (a) Expressiveness. Language-based IDs preserve rich semantic and structural information from open-world settings. (b) Transferability. Like human lexicons, the vocabulary is grounded in natural language, providing a shared foundation that ensures cross-graph interpretability and transferability. (c) Scalability. Any node, whether previously seen or not, can dynamically generate its language-based ID, keeping compatibility with existing nodes while preventing token explosion. We analyze its coverage and uniqueness in Section 5.6.
4.2.2. Multi-Instruction Fine-Tuning.
@@
As shown in Figure 4, we employ a multi-instruction fine-tuning framework to incorporate diverse graphs and tasks (Chung et al., 2024; Wei et al., 2022). We index nodes from the vocabulary and embed their language-based IDs into task-oriented templates to construct completed instructions $\mathcal{T}$:
(6) $\mathcal{T}\leftarrow\text{Prompt}_{\text{template}}\left(T,G\mid\mathcal{F}\right),$
@@
which include language-based IDs of the central nodes, along with its corresponding local structure from $G$ depending on specific graph-centric tasks (e.g., node classification or link prediction). As these instructions are composed entirely of natural language tokens, they are fully readable and coherent. All graph-centric tasks are then reformulated in a text-to-text format, allowing the LLM to be fine-tuned within a unified instruction learning paradigm (Mishra et al., 2022). Given a target sequence $Y$, the loss function can be computed as:
(7) $\mathcal{L}$ $=-\sum_{j=1}^{|Y|}\log\text{Pr}\left(y_{j}\mid\mathcal{T},Y_{<j}\right),$
@@
where $\text{Pr}\left(y_{j}\mid\mathcal{T},Y_{<j}\right)$ denotes the probability of the $j$-th token $Y_{j}$ in the output sequence $Y$, conditioned on the instruction $\mathcal{T}$ and all previous tokens $Y_{<j}=(y_{1},y_{2},\dots,y_{j-1})$. To support broad generalization, we fine-tune LLM backbones such as T5 (Raffel et al., 2020), FLAN (Wei et al., 2022), and Llama (Touvron et al., 2023) to co-train a unified model across diverse graphs and tasks. As a result, the model acquires open-world knowledge and generalizes effectively to previously unseen graphs and tasks.
4.2.3. Constrained Decoding via Prefix Tree Search.
@@
To mitigate LLM hallucination, we introduce a constrained decoding strategy based on a prefix tree search to regulate LLM outputs  (Cao et al., 2021; Tan et al., 2024). Taking generative link prediction as an example, we craft a prefix tree from language-based IDs of all candidate nodes, where each tree node corresponds to a natural language token and each root-to-leaf path uniquely identifies a valid node ID. During autoregressive generation, token selection is constrained to a valid path within the prefix tree, ensuring that predictions align with actual graph nodes and eliminating hallucinations. This effectiveness can be attributed to the discrete nature of language-based IDs, further highlighting the flexibility of our proposed graph vocabulary.
4.3. Efficiency Analysis.
4.3.1. Graph Understanding Module.
Although PromptGFM incurs an initial API expense during the graph understanding phase, this cost is a one-time overhead and remains manageable with modern infrastructure. Each graph is processed only once to generate language-based IDs, which are cached and reused across the inference pipeline. Consequently, when new graphs are introduced, existing IDs do not need to be regenerated via repeated API calls; instead, additional LLM fine-tuning over these natural language tokens is sufficient to capture updated features.
4.3.2. Graph Inference Module.
With IDs cached, both fine-tuning and decoding operate entirely on tokens of these IDs. At inference time, the complexity per decoding step is $O(L)$, where $L$ is the average ID length. Prefix tree search constrains the candidate space to graph-specific vocabulary size ($N$), the overall decoding complexity becomes $O(N\times L)$. These properties make the computational overhead manageable, even in the large-scale graph setting.
Table 1. Evaluation results ($\%$) on node classification accuracy ($\uparrow$). We highlight the methods with the first and second best performances, and report the statistical significance over the strongest baseline using a two-tailed $t$-test with $p<0.05$.
@@
pp-value 4e-3 2.73e-8 3.21e-7 4.48e-10 4.85e-7 3.62e-2 9.93e-4
5. Experiments
In this section, we conduct extensive experiments to address the following research questions (RQs): RQ1: How does PromptGFM perform on node classification and link prediction tasks? RQ2: How to transfer across diverse graphs and tasks as a general-purpose GFM? RQ3: How does PromptGFM accommodate more pretraining tasks? RQ4: What factors influence the replication of prompt-based GNN? RQ5: What is the computational efficiency in graph understanding and inference modules?
5.1. Experimental Setup
5.1.1. Datasets.
We present following datasets from three domains: (1) Computer Science: Cora (McCallum et al., 2000), Citeseer (Giles et al., 1998), Ogbn-arxiv (Hu et al., 2020), and WikiCS  (Mernyei and Cangea, 2020); (2) E-commerce: Photo and History (Ni et al., 2019; Yan et al., 2023); (3) Biomedical: PubMed (Sen et al., 2008). Details can be found in Appendix A.1.
5.1.2. Baselines.
We make comprehensive comparisons with existing methods in four categories: (1) Graph-agnostic methods. We consider the MLP model without graph structure. (2) GNN-based methods. We employ four fundamental GNN models: GCN  (Kipf and Welling, 2017), GAT (Velickovic et al., 2017), GraphSAGE (Hamilton et al., 2017), and ReVGNN (Li et al., 2021). We also explore SGFormer (Wu et al., 2023) and NodeFormer (Wu et al., 2022), which leverage Transformer architectures to model graph data. (3) LM-only methods. We utilize BERT (Devlin et al., 2019), Sentence-BERT  (Reimers and Gurevych, 2019), and Llama3-8B (Touvron et al., 2023) to address downstream tasks. (4) GNN-LLM Integration methods. Following the proposed taxonomy, GraphPrompter  (Liu et al., 2024b) is an instance of using GNNs to enhance LLMs. OFA  (Liu et al., 2024a) and ENGINE  (Zhu et al., 2024) are examples of leveraging LLMs for GNNs. Besides, we incorporate LLaGA  (Chen et al., 2024c) as an attempt of implementing LLM as GNN. Details can be found in Appendix A.2.
5.1.3. Reproduction Settings.
We conduct all experiments in PyTorch on four NVIDIA RTX A6000 GPUs. The graph understanding module uses OpenAI’s GPT-4o mini while we fine-tune a Flan-T5 or Llama3-8B in the graph inference module. We employ 10-fold cross-validation and compute average results with standard deviations. For evaluation, we use accuracy and Macro-F1 (M-F1) for node classification, and accuracy and HR@1 for link prediction. To align with the TAG setting, BERT is used for node initialization across all applicable baselines. More details are provided in Appendix A.3.
Table 2. Accuracy ($\%$) on discriminative link prediction.
Method Discriminative: Accuracy ($\uparrow$)
@@ 5.1. Experimental Setup
Table 3. HR@1 ($\%$) on generative link prediction.
Method Generative: HR@1 ($\uparrow$)
@@
5.2. Performance Comparison (RQ1)
5.2.1. Main Results.
We train each model independently from scratch on a single graph and compare their performance. Tables 1 and 2 summarize the results for node classification and link prediction accuracy, respectively. PromptGFM achieves substantial improvements over state-of-the-art models, leading to the following observations. (a) Graph-based models consistently outperform graph-agnostic ones, highlighting the importance of structural information. (b) Our method surpasses OFA, ENGINE, and GraphPrompter on node classification, suggesting the limitations of their decoupled architectures in both GNN-for-LLM and LLM-for-GNN paradigms. (c) PromptGFM also surpasses LLaGA, which relies on templates to implicitly understand graph semantics and structures. These results indicate that heuristic prompting alone is insufficient to capture high-order dependencies without a true GNN mechanism. In contrast, PromptGFM effectively replicates message passing through a prompt-based GNN, demonstrating the feasibility of using LLMs as GNNs and establishing a new paradigm for graph-text alignment. We provide further analysis in Section 5.5.
5.2.2. Generative Link Prediction.
Following a transductive setting, we partition the graph by links, construct an input graph from the training edges, and predict missing connections for test nodes, treating existing nodes as candidate targets. From Table 3, PromptGFM consistently outperforms traditional GNN models on generative link prediction. Existing GNN-LLM approaches, however, fail in this task: their reliance on OOV token embeddings prevents direct mapping from LLM outputs to specific nodes, often leading to hallucinated predictions. In contrast, our method represents nodes as finite token sequences and employs constrained decoding through prefix tree search, which regulates LLM outputs. This further underscores the significance of our graph vocabulary and demonstrates the flexibility of PromptGFM in generative scenarios.
Table 4. Intra-domain cross-graph transferability to Citeseer within the computer science domain.
@@ 5.2. Performance Comparison (RQ1)
Table 5. Inter-domain cross-graph transferability from e-commerce to biomedical domain.
@@
Table 6. Inter-domain cross-graph transferability from computer science to biomedical domain.
@@
5.3. Cross-graph and Cross-task Transfer (RQ2)
We investigate cross-graph transferability in both intra-domain and inter-domain scenarios. Following  (Chen et al., 2024a), we explore two settings: (a) Pre-training, where models are evaluated on entirely unseen datasets without prior exposure during training, and (b) Co-training, where a portion of the target dataset is available and jointly used with auxiliary datasets for training. These two dimensions yield four distinct transfer paradigms, enabling a comprehensive assessment of cross-graph transferability as follows.
5.3.1. Intra-domain Cross-graph Transferability.
As illustrated in Table 4, all settings surpass direct inference using an off-the-shelf LLM. Notably, incorporating Arxiv yields greater gains than Cora, suggesting that larger and richer datasets enhance transferability more effectively. In the pre-training setting, substantial improvements over direct inference demonstrate strong zero-shot transferability within the domain. A similar, albeit smaller, trend is observed in the co-training setting, further validating the effectiveness of intra-domain knowledge transfer. These findings underscore the potential to collect more graph data and build more knowledgeable and transferable GFMs.
5.3.2. Inter-domain Cross-graph Transferability.
We report cross-domain transfer results from different domains to the biomedical domain in Tables 5 and 6, respectively. In the pre-training setting, while incorporating a single external graph always enhances performance, adding multiple source graphs may lead to negative transfer. More critically, in the co-training setting under supervised learning, performance consistently deteriorates as more cross-domain data is introduced. This decline mainly stems from domain shifts, including structural differences, semantic drift in text, and mismatched label spaces  (Chen et al., 2024a; Zhao et al., 2024). In contrast, PromptGFM adopts a unified language-based vocabulary and pure language interface, which enables natural cross-domain transferability and avoids token collision in principle. While this remains a broader challenge for the community, we provide additional discussion and supporting experiments in Appendix C.2.
Table 7. Cross-task generalization between node classification (NC) and link prediction LP. ST-X refers to supervised training on the task X from scratch.
Task Setting Cora Citeseer PubMed
NC zero-shot 18.54 27.64 39.12